OpenStudy (anonymous):

A student is trying to solve the system of two equations given below: Equation P: a + b = 6 Equation Q: 4a + 2b = 19 Which of the following steps can be used to eliminate the a term? A) -1(4a+2b=19) B) -4(4a+2b=19) C) -4(a+b=6) D) 4 (a+b=6)

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

it should be D

OpenStudy (anonymous):

a + b = 6 4a + 2b = 19........................... for you to eleminate a, multiply the 1st equation by 4 and the second equation by 1 .......

OpenStudy (anonymous):

thanks @GIL.ojei

OpenStudy (anonymous):

4 x ( a + b = 6) 1 x ( 4a + 2b = 19)

OpenStudy (anonymous):

d right ? @GIL.ojei

OpenStudy (anonymous):

yees

OpenStudy (133):

No it is not D